Need help please. This question makes no sense, and I can't find an example that explains what needs to be done to solve it.
n(U) = 100, n(A) = 24, n(B) = 60, and n(A ∩ B) = 10. Find n(A' ∩ B).
Thank you!

Answer by ikleyn(52803)

.
The set  (A' ∩ B)  is the set of those elements of B, that do not belong to A.


So  (A' ∩ B)  is simply the difference of two subsets  B  and  (A ∩ B) :


    (A' ∩ B) = B \ (A ∩ B),      (*)


where the symbol  " \ "  symbolizes subtraction of subsets.


From (*),  n(A' ∩ B) = n(B) - n(A ∩ B) = 60 - 10 = 50.    ANSWER

Solved.